The length of a rectangleis 6m more than 4 times the width. The perimeter is 72m. Find the length and the width
Answer by Maths68(1474)

Perimeter of a Rectangle = 72 meters
Length = l=4w+6
Width = w
Perimeter of a Rectangle = length + length + width + width
Perimeter of a Rectangle = 2l + 2w
72 = 2(4w+6) + 2w
72 = 8w+12 + 2w
72-12 = 10w
60=10w
60/10=w
w=6
Length = l=4w+6 (Given)
Length = l=4(6)+6
Length = l=24+6=30
Dimensions of the rectagle
Length = 30m, Width=6m, Perimeter=72m